# Beyond Hawking evaporation of black holes formed by dark matter in compact stars

### v1

@@ I Introduction
Figure 1: Limits on the dark matter-neutron scattering cross section as a function of dark matter mass from the survival of the neutron star PSR J0437--4715; the top x-axis shows the initial mass of dark matter-seeded black holes corresponding to the bottom x-axis, as obtained from Eqs. (4) and (5). The red shaded region shows our extension of previous limits, namely, the blue shaded region, when suppressed evaporation of these black holes via the memory burden effect is taken into account. The gray vertical line shows the dark matter mass above which the neutron star will always survive due to the black hole Bondi-accreting stellar material on a timescale smaller than the age of the star. Also shown for comparison are recent constraints from a direct search at LZ Aalbers et al. (2024) obtained by linearly extrapolating their exclusion cross sections at high dark matter masses. See Sec. III for more details.
@@
The suppression of a black hole’s evaporation at the end stages of its decay via Hawking radiation is predicted by several effects, such as when black holes are magnetically charged Stojkovic and Freese (2005); Maldacena (2021); Estes et al. (2023); Banerjee et al. (2024), attain extremality near the Planck mass Chen and Adler (2003); Alexeyev et al. (2002); Chen (2005); Nozari and Mehdipour (2005); Lehmann et al. (2019); Profumo (2024), or quasi-extremality de Freitas Pacheco et al. (2023), or form massive remnants Giddings (1992). Recently, Refs. Dvali (2018); Dvali et al. (2020, 2024) proposed a mechanism by which the breakdown of the Hawking treatment occurs at much earlier times, at the latest when the black hole has shed half its mass, i.e. around the Page time. The basic principle is that the very large entropy of a black hole corresponds to a very high capacity for storage of information, which is possible through gapless degrees of freedom that can be arranged at a low energy cost. The initial stages of a black hole’s evaporation could then be modeled semi-classically as in the Hawking picture, however, the information remaining within the black hole produces a backreaction on the black hole itself. The backreaction acts to increase the gaps in the energy levels, and quantum effects can no longer be neglected in modeling the radiation, which is now suppressed. Therefore the evaporation is no longer self-similar until the black hole shrinks to the extremal scale – now it does have memory. This “memory burden effect” is generic to any decaying system with high information storage capacity, e.g. saturons Dvali et al. (2024). It provides an explicit mechanism for release of information from a black hole after the Page time, thereby resolving the information paradox, and gives it a quantum hair that is not constrained by classical no-hair theorems. The phenomenological consequences of slow evaporation of black holes after half-decay have recently been used to derive novel limits on a number of scenarios, including primordial black holes as dark matter Alexandre et al. (2024); Thoss et al. (2024); Balaji et al. (2024); Dvali et al. (2024); Barman et al. (2024a); Bhaumik et al. (2024); Barman et al. (2024b); Kohri et al. (2024); Zantedeschi and Visinelli (2024); Chianese et al. (2024); Barker et al. (2024); Borah and Das (2024); Loc (2024). In this work, we use this framework for the first time to constrain scenarios involving particle dark matter.
@@ II.1 Capture, thermalization, collapse, black hole formation
where $\rho_{\chi}$ and $\bar{v}_{\rm gal}$ are the density and average speed of the ambient dark matter, and the $A^{2}$ factor accounts for inefficient momentum transfers at large $m_{\chi}$ Bramante et al. (2017); Dasgupta et al. (2019).
@@
$r_{\rm th}\simeq 0.03\,{\rm cm}\,\left(\frac{10^{9}~{}\rm GeV}{m_{\chi}}\right)^{1/2}\left(\frac{T_{\rm NS}}{2\times 10^{6}\,\rm K}\right)^{1/2}\\ \times\left(\frac{1.4\times 10^{15}\,\rm g/cm^{3}}{\rho_{\rm NSc}}\right)^{1/2}\,,$ (3)
@@
$M_{\chi}^{\rm self}=1.5\times 10^{35}~{}{\rm GeV}\left(\frac{10^{9}\,\rm GeV}{m_{\chi}}\right)^{3/2}\left(\frac{T_{\rm NSc}}{2\times 10^{6}\,{\rm K}}\right)^{3/2}\\ \times\left(\frac{1.4\times 10^{15}\,\rm g/cm^{3}}{\rho_{\rm NSc}}\right)^{1/2}\,.$ (4)
@@
where for the spin-0 dark matter we have assumed no self-interactions. A black hole is only formed after the mass captured over a time $t$, which is $m_{\chi}C_{\chi}t$, reaches max[$M_{\chi}^{\rm self}$, $M_{\chi}^{\rm Ch}$]. For our parametric range of interest, black hole formation for spin-0 dark matter is set by the self-gravitation criterion, and for spin-1/2 dark matter by the Chandrasekhar criterion.
@@
where $P(M_{\text{BH}})$ is the Page factor accounting for gray-body corrections of the Hawking evaporation spectrum, as well as the number of Standard Model species emitted from an evaporating black hole Page (1976); MacGibbon (1991); Ray (2023). As the black hole temperature $\propto 1/M_{\rm BH}$, in the course of evaporation the number of particle species emitted, hence the Page factor, increases. For $M_{\rm BH}>10^{17}$ g, $P(M_{\rm BH})=1/1135\pi$ with only photons and neutrinos emitted. For $M_{\rm BH}\ll 10^{10}$ g, $P(M_{\rm BH})=1/74\pi$ with all Standard Model states emitted. We use Ref. MacGibbon (1991) to obtain the semi-analytical form of the Page factor, which agrees reasonably well with BlackHawk Arbey and Auffinger (2019). In Eq. (6) we have conservatively omitted a growth rate term coming from ongoing dark matter accretion on the neutron star; it is generally negligible compared to the nucleon Bondi accretion rate, and the accretion of dark matter on the black hole is model-dependent Garani et al. (2019).
@@ II.2 The effect of memory burden
where $S(M_{\rm BH})=4\pi GM^{2}_{\rm BH}$ denotes the entropy of the black hole, and $k$ is restricted to positive integers since the lifetime of the black hole is an analytic function of the entropy Dvali et al. (2024). Note that $k=0$ corresponds to the standard Hawking scenario. Since the entropy is a very large number the $S(M_{\rm BH})^{k}$ term has the effect of suppressing evaporation. The memory burden effect may also precipitate a classical instability that triggers rapid decay of the black hole Dvali et al. (2020), but we do not consider this possibility.
@@ III Results
In Fig. 1 we show our limits for spin-0 (left panel) and spin-1/2 (right panel) dark matter in the plane of spin-independent nucleon scattering cross section and dark matter mass. In the top x-axis we mark the initial mass of the black hole corresponding to the $m_{\chi}$ in the bottom x-axis, as obtained from Eqs. (4) and (5). The shaded regions are excluded by the continued existence of PSR J0437--4715. We also show limits from a recent search at the underground experiment LZ  Aalbers et al. (2024), which are orders of magnitude weaker than our limits in the relevant mass range. This is because our neutron star provides much larger target exposure to incident dark matter flux.
@@
We revise former limits by extending them into the red shaded region, where suppressed evaporation allows lighter black holes to consume the neutron star. We have taken the fraction of mass lost in the Hawking stage of evaporation $f_{\rm loss}$ as 1/2, and taken the suppression exponent $k$ in Eq. (8) as 1. These choices already push the limit on $m_{\chi}$ to within a factor of 2 of the Bondi accretion limit in Eqs. (9) and (10). Refs. Dvali (2016); Michel and Zell (2023) have shown that quantum backreaction could set in at an earlier time and suppress black hole evaporation. In this light, we find that for smaller values of $f_{\rm loss}$, despite a longer period of suppressed evaporation from memory burden, our limits do not improve by much because we rapidly run into the Bondi accretion limit in Eqs. (9) and (10), i.e. the gray vertical line in Fig. 1. We also find that setting $k$ to integer values $>1$ has no effect on the limits in Fig. 1. This is because of the great entropy-dependent suppression in the evaporation rate of Eq. (8), which practically switches off black hole decay for any integer $k$ in our parameter range of interest.

### v2

@@ I Introduction
Figure 1: Limits on the dark matter-neutron scattering cross section as a function of dark matter mass from the survival of the neutron star PSR J0437--4715; the top x-axis shows the initial mass of dark matter-seeded black holes corresponding to the bottom x-axis, as obtained from Eqs. (4) and (5). The red shaded region shows our extension of previous limits, namely, the blue shaded region, when suppressed evaporation of these black holes via the memory burden effect is taken into account. The gray vertical line shows the dark matter mass above which the neutron star will always survive due to the black hole Bondi-accreting stellar material on a timescale smaller than the age of the star. Also shown for comparison are recent constraints from a direct search at LZ Aalbers et al. (2024) and the neutrino fog  O’Hare (2021), obtained by linearly extrapolating exclusion cross sections at high dark matter masses. We have assumed here that half the mass of the black hole is lost before the memory burden phase, and that the exponent $k=1$ in Eq. (8), but altering these choices have little effect on our limits.
@@
The suppression of a black hole’s evaporation at the end stages of its decay via Hawking radiation is predicted by several effects, such as when black holes are magnetically charged Stojkovic and Freese (2005); Maldacena (2021); Estes et al. (2023); Banerjee et al. (2024), attain extremality near the Planck mass Chen and Adler (2003); Alexeyev et al. (2002); Chen (2005); Nozari and Mehdipour (2005); Lehmann et al. (2019); Profumo (2024), or quasi-extremality de Freitas Pacheco et al. (2023), or form massive remnants Giddings (1992). Recently, Refs. Dvali (2018); Dvali et al. (2020, 2024) proposed a mechanism by which the breakdown of the Hawking treatment occurs at much earlier times, at the latest when the black hole has shed half its mass, i.e. around the Page time. The basic principle is that the very large entropy of a black hole corresponds to a very high capacity for storage of information, which is possible through gapless degrees of freedom that can be arranged at a low energy cost. The initial stages of a black hole’s evaporation could then be modeled semi-classically as in the Hawking picture, however, the information remaining within the black hole produces a backreaction on the black hole itself. The backreaction acts to increase the gaps in the energy levels, and quantum effects can no longer be neglected in modeling the radiation, which is now suppressed. Therefore the evaporation is no longer self-similar until the black hole shrinks to the extremal scale – now it does have memory. This “memory burden effect” is generic to any decaying system with high information storage capacity, e.g. saturons Dvali et al. (2024). It provides an explicit mechanism for release of information from a black hole after the Page time, thereby resolving the information paradox, and gives it a quantum hair that is not constrained by classical no-hair theorems. The phenomenological consequences of slow evaporation of black holes after half-decay have recently been used to derive novel limits on a number of scenarios, including primordial black holes as dark matter Alexandre et al. (2024); Thoss et al. (2024); Balaji et al. (2024); Haque et al. (2024); Dvali et al. (2024); Barman et al. (2024a); Bhaumik et al. (2024); Barman et al. (2024b); Kohri et al. (2024); Zantedeschi and Visinelli (2024); Chianese et al. (2024); Barker et al. (2024); Borah and Das (2024); Loc (2024). In this work, we use this framework for the first time to constrain scenarios involving particle dark matter.
@@ II.1 Capture, thermalization, collapse, black hole formation
where $\rho_{\chi}$ and $\bar{v}_{\rm gal}$ are normalized to the ambient density and average Maxwell-Boltzmann speed of dark matter particles, and the $A^{2}$ factor accounts for inefficient momentum transfers at large $m_{\chi}$ Bramante et al. (2017); Dasgupta et al. (2019).
@@
$r_{\rm th}\simeq 0.03\,{\rm cm}\,\left(\frac{10^{9}~{}\rm GeV}{m_{\chi}}\right)^{1/2}\left(\frac{T_{\rm NS}}{2.1\times 10^{6}\,\rm K}\right)^{1/2}\\ \times\left(\frac{1.4\times 10^{15}\,\rm g/cm^{3}}{\rho_{\rm NSc}}\right)^{1/2}\,,$ (3)
@@
$M_{\chi}^{\rm self}=1.5\times 10^{35}~{}{\rm GeV}\left(\frac{10^{9}\,\rm GeV}{m_{\chi}}\right)^{3/2}\left(\frac{T_{\rm NSc}}{2.1\times 10^{6}\,{\rm K}}\right)^{3/2}\\ \times\left(\frac{1.4\times 10^{15}\,\rm g/cm^{3}}{\rho_{\rm NSc}}\right)^{1/2}\,.$ (4)
@@
where for the spin-0 dark matter we have assumed no self-interactions. A black hole is only formed after the mass captured over a time $t$, which is $m_{\chi}C_{\chi}t$, reaches max[$M_{\chi}^{\rm self}$, $M_{\chi}^{\rm Ch}$]. For our parametric range of interest, black hole formation for spin-0 dark matter is set by the self-gravitation criterion, and for spin-1/2 dark matter by the Chandrasekhar criterion. In principle a small fraction of the nascent mass of the black hole may be emitted as gravitational radiation at the time of black hole formation East and Lehner (2019); we assume for simplicity that this is avoided by spherically symmetric collapse of dark matter.
@@
where $P(M_{\text{BH}})$ is the Page factor accounting for gray-body corrections of the Hawking evaporation spectrum, as well as the number of Standard Model species emitted from an evaporating black hole Page (1976); MacGibbon (1991); Ray (2023). As the black hole temperature $\propto 1/M_{\rm BH}$, in the course of evaporation the number of particle species emitted, hence the Page factor, increases. For $M_{\rm BH}>10^{17}$ g, $P(M_{\rm BH})=1/1135\pi$ with only photons and neutrinos emitted. For $M_{\rm BH}\ll 10^{10}$ g, $P(M_{\rm BH})=1/74\pi$ with all Standard Model states emitted. We use Ref. MacGibbon (1991) to obtain the semi-analytical form of the Page factor, which agrees reasonably well with BlackHawk Arbey and Auffinger (2019). In Eq. (6) we have conservatively omitted a growth rate term coming from ongoing dark matter accretion on the neutron star; it is generally negligible compared to the nucleon Bondi accretion rate, and the accretion of dark matter on the black hole is model-dependent Garani et al. (2019). We also neglect possible emission of non-standard states (such as $\chi$) from the nascent black hole, which can enhance evaporation rates and weaken our constraints.
@@ II.2 The effect of memory burden
where $S(M_{\rm BH})=4\pi GM^{2}_{\rm BH}$ denotes the entropy of the black hole, and $k$ is restricted to positive integers in order to keep the lifetime of the black hole an analytic function of the entropy Dvali et al. (2024). Note that $k=0$ corresponds to the standard Hawking scenario. In this work we will set $k=1$ to obtain results, but we find that these are practically unchanged for any $k\geq 1$. This is because, in our parameter range of interest, the enormous entropy of the black hole (e.g., $S(10^{12}~{}{\rm g})\sim 10^{34}$) practically switches off its evaporation in the memory-burdened phase. The memory burden effect may also precipitate a classical instability that triggers rapid decay of the black hole Dvali et al. (2020), but we do not consider this possibility, and can be addressed in a future work.
@@ III Results
In Fig. 1 we show our limits for spin-0 (left panel) and spin-1/2 (right panel) dark matter in the plane of spin-independent nucleon scattering cross section and dark matter mass. In the top x-axis we mark the initial mass of the black hole corresponding to the $m_{\chi}$ in the bottom x-axis, as obtained from Eqs. (4) and (5). The shaded regions are excluded by the continued existence of PSR J0437--4715. We also show limits from a recent search at the underground experiment LZ Aalbers et al. (2024). Also shown are “neutrino fog” cross sections O’Hare (2021) below which dark matter direct detection is expected to be challenging due to an irreducible background from atmospheric neutrinos. These limits are orders of magnitude weaker than our limits in the relevant mass range. This is simply because celestial objects, such as, neutron stars provide significantly larger target exposure to the incident dark matter flux. We note here, though, that direct detection limits are agnostic to the annihilation properties of dark matter, whereas we require negligible annihilations for black hole formation.
@@
We revise former limits by extending them into the red shaded region, where suppressed evaporation allows lighter black holes to consume the neutron star. We have taken the fraction of mass lost in the Hawking stage of evaporation $f_{\rm loss}$ as 1/2, and taken the suppression exponent $k$ in Eq. (8) as 1. These choices already push the limit on $m_{\chi}$ to within a factor of 2 of the Bondi accretion limit in Eqs. (9) and (10). Refs. Dvali (2016); Michel and Zell (2023) have shown that quantum backreaction could set in at an earlier time and suppress black hole evaporation. In this light, we find that for smaller values of $f_{\rm loss}$, despite a longer period of suppressed evaporation from memory burden, our limits do not improve by much because we rapidly run into the Bondi accretion limit in Eqs. (9) and (10), i.e. the gray vertical line in Fig. 1. As mentioned before, we also find that setting $k$ to integer values $>1$ has no effect on the limits in Fig. 1 due to the entropy-dependent suppression in the evaporation rate in Eq. (8).